OpenStudy (anonymous):

The polynomial 32x^2+12x-57 has two real roots. A. True B. False

OpenStudy (anonymous):

A

OpenStudy (lgbasallote):

we calculate for b^2 - 4ac in here b = -12 a = 32 c = -57 \[\large b^2 - 4ac = (-12)^2 - 4(32)(-57)\] \[\large b^2 - 4ac = 144 + 4(32)(57)\] from here we can plainly see that the answer will be positive so there are 2 real roots

OpenStudy (anonymous):

the degree of the equation (highest exponent) is the number of real roots

OpenStudy (lgbasallote):

@DreaWolfChild the degree of the exponent is the number of roots..but it does not determine whether the roots are real or imaginary. to get that you have to solve for the discriminant :)